### v1

# The open driven two-level system at conical intersections of quasienergies

Abstract
We study the stationary state of an ac-driven two-level system under particle exchange with a fermionic environment. A particular question addressed is whether there exist limits in which the populations of the Floquet states are determined by their quasienergies or their mean energies, respectively. The focus lies on parameters in the vicinity of conical intersections of quasienergies, because there the two kinds of energies behave rather differently, such that the characteristics of the two intuitive limits are most pronounced. A main finding is a crossover from a Floquet-Gibbs state at low temperatures to a mean-energy dominated state at intermediate temperatures. Analytical estimates are confirmed by numerical calculations.
I Introduction
Quantum systems at thermal equilibrium are usually assumed to be in a Gibbs states, depending on whether one allows for particle exchange, it will be the canonical or the grand canonical ensemble. For a more profound description, one may employ system-bath models that allow for the exchange of energy Weiss (2012) or of particles Bruus and Flensberg (2004) with an environment to obtain the dynamics and the stationary solution of the reduced system density operator. Usually at least for very weak coupling, the latter eventually becomes a Gibbs state.Landi, Poletti, and Schaller (2022) For driven systems, the situation is more involved, because no such formal expression for the stationary density operator is known. Hence, for its computation, system-bath models are indispensable. Notably, then even in the weak-coupling limit, the stationary state may qualitatively depend on details of the system-bath interaction.Ferrón, Domínguez, and Sánchez (2012); Blattmann, Hänggi, and Kohler (2015) Moreover, genuine non-equilibrium effects such as pumping and rectification of heat Segal and Nitzan (2005, 2006) and charge Lehmann et al. (2002, 2003) emerge. Such non-equilibrium effects are particularly important for Floquet engineering which is the design of effective static Hamiltonians by ac fields.Grossmann et al. (1991); Großmann and Hänggi (1992); Holthaus (1992); Eckardt (2017) In exceptional cases of driven dissipative systems, there emerge Floquet-Gibbs states, i.e., canonical states in which the eigenenergies are formally replaced by quasienergies.Thingna, Wang, and Hänggi (2012); Shirai, Mori, and Miyashita (2015); Shirai et al. (2016); Iwahori and Kawakami (2016); Mori (2023) There are also situations in which, by contrast, the mean energies of the Floquet states determine the populations.Kohler et al. (1998); Ketzmerick and Wustmann (2010); Kohler (2024)
In practice none of these limits will be perfectly realized. Nevertheless one may observe a tendency towards the one or the other. To explore this behavior, the vicinity of conical intersections of quasienergies turned out rather to be interesting.Kohler (2024) Such intersections emerge when a spatio-temporal symmetry of a driven quantum system such as generalized parity Peres (1991) allows the exact crossing of quasienergies. A weak perturbation may break this symmetry, so that the former crossing turns into an accidental degeneracy in a two-dimensional parameter space. A most interesting feature of canonical intersections in driven systems is that along lines with constant quasienergies, the mean energies of the Floquet states are interchanged.Kohler (2024) Therefore in this regime, the roles played by these energies will be most pronounced. For example, there may emerge discontinuities of the populations Engelhardt, Platero, and Cao (2019) or a crossover from a mean-energy state to a more exceptional Floquet-Gibbs state.Kohler (2024)
In this work, we explore how the results of Ref. Kohler, 2024 for the driven dissipative two-level system are affected by replacing the bosonic heat bath with electron reservoirs. To stay close to that situation, electron-electron interaction and the spin degree of freedom are ignored. A suitable description of such systems is a Floquet-based master equation for the single-particle density matrix. It was derived in the context of molecular wires to study ratchet effects Lehmann et al. (2002) and current rectification by ac fields. Lehmann et al. (2003)
This work is structured as follows. In Sec. II, we introduce conical intersections of quasienergies as a consequence of a spatio-temporal symmetry, the coupling to a fermionic particle reservoir, and a master equation formalism for the single-particle density operator based on Floquet theory. Section III is devoted to approximative solutions that lead to conjectures for the stationary state which are confirmed by numerical calculations. Conclusions are drawn in Sec. IV, while the properties of Floquet states close to conical intersections are summarized in the appendix.
II Open driven two-level system
II.1 Driven two-level system and conical intersections
Figure 1: Driven two-level system with tunnel coupling $\Delta$ and ac-modulated onsite energies. The wire-lead tunnel couplings $\Gamma_{L,R}$ enable particle exchange with the respective site.
Figure 2: (a) Quasienergies (solid lines) and mean energies (dashed) for driving frequency $\Omega=10\Delta$ and the rather small detuning $\varepsilon=0.03\Delta$ as function of the driving amplitude. For $\varepsilon=0$, owing to generalized parity the avoided crossing would be exact. (b) Quasienergies as a function of the detuning and the driving amplitude in the vicinity of the first quasienergy crossing observed in panel (a) close to $\varepsilon=0$ and $A_{0}\approx 2.4\Omega$. (c) Mean energy of the Floquet state $|\phi_{0}(t)\rangle$ with—for the Brillouin zone centered at zero—lower quasienergy.
We consider a two-level system with tunable onsite energies and tunnel coupling to two leads as is sketched in Fig. 1. When occupied with a single electron, the Hamiltonian of the central system in pseudo-spin notation reads
$H(t)=\frac{\Delta}{2}\sigma_{x}+\frac{1}{2}[\varepsilon+A\cos(\Omega t)]\sigma_{z}\,,$ (1)
where $\sigma_{x,z}$ are the usual Pauli matrices. The two sites are tunnel coupled with strength $\Delta$ and possess a static detuning $\varepsilon$. The ac driving enters via a sinusoidal additional detuning with amplitude $A$ and frequency $\Omega$. We use units with $\hbar=1$.
A suitable tool for treating periodically time-dependent quantum systems is Floquet theory. Its cornerstone is discrete time translation invariance, from which follows that the Schrödinger equation has a complete set of solutions of the form $e^{-iqt}|\phi(t)\rangle$ with quasienergy $q$ and Floquet state $|\phi(t)\rangle=|\phi(t+2\pi/\Omega)\rangle$. Shirley (1965); Sambe (1973); Hänggi (1998) The Floquet states and their quasienergies are eigensolutions of the operator $\mathcal{H}=H(t)-i\partial/\partial t$ in Sambe space, which is the direct product of Hilbert space and the space of $2\pi/\Omega$-periodic functions.Sambe (1973) Quasienergies correspond to the phase that a Floquet state acquires during one driving period. An important property of a Floquet state is its mean energy $E$, i.e, its energy expectation value averaged over one driving period. It relates to the quasienergy via a geometric phase,Aharonov and Anandan (1987); Moore (1990)
$E=q+\langle\!\langle\phi(t)|i\partial_{t}|\phi(t)\rangle\!\rangle=q+\sum_{k}k\Omega\langle\phi_{k}|\phi_{k}\rangle,$ (2)
where the outer angles denote time average, while $|\phi_{k}\rangle$ is the $k$th Fourier coefficient of the Floquet state.
When $|\phi(t)\rangle$ is a Floquet state with quasienergy $q$, then for any integer $n$, $e^{-in\Omega t}|\phi(t)\rangle$ is a physically equivalent Floquet state with quasienergy $q+n\Omega$. All equivalent states correspond to the same solution of the Schrödinger equation. Therefore, it is sufficient to consider a particular one. Nevertheless in the present context, there exists a unique meaningful choice, namely the one for which upon adiabatically reducing the driving amplitude to zero, $q$ and $E$ become equal. Only then we can expect the emergence of a meaningful Floquet-Gibbs state. This choice can be ensured by considering large driving frequencies and choosing the Brillouin zone such that it comprises the spectrum of the undriven Hamiltonian.Shirai et al. (2016) This implicitly fixes the ordering of the Floquet states such that we can attribute the “ground state index 00” to the state with lower quasienergy.
For $\varepsilon=0$, both the Hamiltonian (1) and the Floquet Hamiltonian $\mathcal{H}$ possess a spatio-temporal $\mathbb{Z}_{2}$ symmetry, namely the generalized parity $G=\sigma_{x}P$, where $P=e^{(\pi/\Omega)\partial_{t}}$ shifts time by half a driving period.Peres (1991) For $2\pi/\Omega$-periodic functions, $G^{-1}=G$ which implies that the eigenvalues of $G$ are ±$\pm 1$. Notably the operators $i\partial_{t}$, $\sigma_{x}$, and $\sigma_{z}\cos(\Omega t)$ are invariant under transformation with $G$ such that $[G,\mathcal{H}]=0$. Therefore, the Floquet states are also eigenstates of $G$ or, in the case of degeneracies, can be chosen as such. Moreover, they can be classified as even or odd, depending on the respective eigenvalue of $G$. It is known that quasienergies from different symmetry classes may form exact crossings.Haake, Gnutzmann, and Kuś (2018)
Since $G\sigma_{z}G^{-1}=-\sigma_{z}$, a non-vanishing detuning $\varepsilon$ breaks the generalized parity of $\mathcal{H}$. As a consequence, exact quasienergy crossings become avoided, as can be witnessed in Fig. 2(a). At avoided crossings, the associated states and, thus, their expectation values are interchanged. This is reflected by the behavior of the mean energies which form exact crossings when quasienergies anti-cross. Moreover, in the parameter space of detuning and driving amplitude, $(\varepsilon,A)$, one finds a degeneracy only at isolated points, such that the quasienergies form conical intersections, see Fig. 2(b). The behavior of the mean energy of the state on the lower cone is sketched in Fig. 2(c). It visualizes the continuous exchange of the mean energies when going around the cone tip along a line with constant quasienergy.Kohler (2024) Accordingly, whenever the mean energies govern dissipation, we can expect a significant parameter dependence of the stationary state. In turn, for a Floquet-Gibbs like behavior, one will observe only minor changes.
Henceforth, we focus on the vicinity of a cone tip at $\varepsilon=0$ and $A=A_{0}$, where the tip position $A_{0}$ depends on $\Delta$ and $\Omega$. In the limit of large driving frequency, the ratio $A_{0}/\Omega$ matches a zero of the zeroth-order Bessel function of the first kind. Here, we do not restrict ourselves to this limit, such that the possible values of $A_{0}$ have to be determined numerically. At the tip, the Floquet states have zero quasienergy and definite generalized parity. We denote these states by $|\varphi_{\pm}(t)\rangle$ with mean energies $E_{\pm}$ and use them states as basis. Then the Floquet Hamiltonian becomes
$\mathcal{H}^{\prime}=\frac{a}{2}\sigma_{z}^{\prime}+\frac{b}{2}\sigma_{x}^{\prime}$ (3)
with the effective parameters
$a={}$ $\frac{A_{0}-A}{A_{0}}\big{(}2E_{-}-\langle\!\langle\varphi_{-}|\sigma_{x}|\varphi_{-}\rangle\!\rangle\big{)},$ (4)
$b={}$ $\varepsilon\langle\!\langle\varphi_{-}|\sigma_{z}|\varphi_{+}\rangle\!\rangle,$ (5)
where the prime refers to the new time-dependent basis. Notice that this constitutes an approximation, because a basis of the full Sambe space contains also all equivalent Floquet states $e^{-in\Omega t}|\varphi_{\pm}(t)\rangle$. The eigenvalues of $\mathcal{H}^{\prime}$ are the quasienergies ±$\pm r/2$ where $r=\sqrt{a^{2}+b^{2}}$. The main ideas of the perturbation theory are summarized in the appendix, while details can be found in Ref. Kohler, 2024.
II.2 Fermionic environment and master equation
To describe the open two-level system, we employ the many-particle version of $H(t)$,
$H_{\text{wire}}(t)=\sum_{n,m}h_{nm}(t)c_{n}^{\dagger}c_{m}\,,$ (6)
where “wire” refers to the central system without the leads. Here, $c_{n}$ is the fermionic annihilation operator of an electron at site $n$, and $h_{nm}$ denotes the matrix elements of the single-particle Hamiltonian (1).
[…]
Following Refs. Lehmann et al., 2002, 2003, we eliminate the leads within second-order perturbation theory to obtain a Bloch-Redfield equation for the reduced density operator of the wire. To benefit from the knowledge acquired from Floquet theory, it is convenient to define the Floquet annihilation operators,
$c_{\alpha}(t)=\sum_{n}\langle\phi_{\alpha}(t)|n\rangle c_{n}$ (7)
which at equal time obey the usual fermionic anti-commutation relations. They allows us to define the Floquet single-particle density operator $R_{\alpha\beta}=\langle c_{\beta}^{\dagger}(t)c_{\alpha}(t)\rangle$ whose equation of motion can be derived from the Bloch-Redfield equation.
For very weak coupling one can apply a rotating-wave approximation which ignores off-diagonal elements of $R$ such that $R_{\alpha\beta}=P_{\alpha}\delta_{\alpha,\beta}$ with the Floquet state populations $P_{\alpha}$. Then in the long-time limit, the populations become practically time-independent and obey Lehmann et al. (2003)
$\sum_{\ell,k}w^{(\ell)}_{\alpha k}P_{\alpha}=\sum_{\ell,k}w^{(\ell)}_{\alpha k}f(q_{\alpha}+k\Omega-\mu_{\ell}).$ (8)
The transitions rates $w^{(\ell)}_{\alpha k}=\Gamma_{\ell}|\langle\ell|\phi_{\alpha k}\rangle|^{2}$ are governed by the sideband resolved overlap of the localized wire state $|\ell\rangle$ with the Floquet state $|\phi_{\alpha}\rangle$. This equation for the $P_{\alpha}$ will be used for analytical considerations, while all numerical results are computed with the full master equation for $R_{\alpha\beta}(t)$.
For the numerical calculations, we first determine the smallest value $A_{0}$ at which the quasienergies for $\varepsilon=0$ cross, i.e., the position of the cone tip. On a circle around the tip with $r\ll\Delta$, the quasienergies are (approximately) constant. To determine the corresponding detuning $\varepsilon$ and amplitude $A$, we parametrize the circle as $a=r\cos\vartheta$ and $b=r\sin\vartheta$, where the angle $\vartheta=0$ stands for a driving amplitude slightly below $A_{0}$, while $\vartheta=\pi/2$ and $23\pi/$ 2, mean $A=A_{0}$ and $\varepsilon\neq 0$. Equations (4) and (5) provide the corresponding values of $\varepsilon$ and $A$.
III Floquet state population
Since owing to the rotating-wave approximation, the populations of the Floquet states in Eq. (8) are uncoupled, one readily obtains the stationary solution
$P_{\alpha}=\frac{\sum_{\ell,k}w^{(\ell)}_{\alpha k}f(q_{\alpha}+k\Omega-\mu_{\ell})}{\sum_{\ell,k}w^{(\ell)}_{\alpha k}}.$ (9)
Sufficiently close to the Fermi surface, it can be approximated by using the Taylor expansion
$f(x)=\frac{1}{2}+\frac{x}{4k_{B}T}\quad\text{for $|x|\lesssim k_{B}T$},$ (10)
while outside this window, $f(x)$ practically equals zero or unity depending on the sign of its argument.
III.1 Equal coupling to both leads
Let us start the discussion with a symmetric situation close to the one with a bosonic bath considered in Ref. Kohler, 2024. We assume that both leads have equal chemical potentials, $\mu_{L}=\mu_{R}=0$, and equal tunnel coupling to the respective lead, $\Gamma_{L}=\Gamma_{R}=\Gamma$. Then we can use the relation $\sum_{\ell}|\ell\rangle\langle\ell|=\mathbf{1}$ to obtain $\sum_{\ell}w_{\alpha k}^{(\ell)}=\Gamma\langle\phi_{\alpha k}|\phi_{\alpha k}\rangle$ and $\sum_{\ell,k}w_{\alpha k}^{(\ell)}=\Gamma$ such that Eq. (9) becomes
$P_{\alpha}=\sum_{k}\langle\phi_{\alpha k}|\phi_{\alpha k}\rangle f(q_{\alpha}+k\Omega).$ (11)
Interestingly, the population is independent of $\Gamma$, despite that the wire-lead couplings $V_{\ell}$ do not commute with the wire Hamiltonian. This result is consistent with the natural expectation for the undriven limit in which the occupation probability of the sidebands becomes $\propto\delta_{k,0}$. Then $P_{\alpha}=f(q_{\alpha})$ with the quasienergy $q_{\alpha}$ being the corresponding eigenenergy of the undriven Hamiltonian.
Expression (11) can be simplified in two limits. First for rather low temperatures, $k_{B}T\lesssim\Omega$, such that the Fermi function for all sideband indices $k<0$ is close unity and practically vanishes for $k>0$, one may assume that the overlaps $\langle\phi_{\alpha k}|\phi_{\alpha k}\rangle$ are independent of the sign of $k$ (which is not exactly fulfilled, but represents a reasonable approximation). Then the normalization of the Floquet states yields $\langle\phi_{\alpha 0}|\phi_{\alpha 0}\rangle+2\sum_{k<0}\langle\phi_{\alpha k}|\phi_{\alpha k}\rangle=0$ such that
$P_{\alpha}=\frac{1}{2}+\langle\phi_{\alpha 0}|\phi_{\alpha 0}\rangle\Big{(}f(q_{\alpha})-\frac{1}{2}\Big{)},$ (12)
i.e., the population is governed by the Fermi function evaluated at the quasienergy. However, it becomes $f(q_{\alpha})$ only if the Floquet state has its full weight in the sideband with $k=0$, which is the case only for $A=0$. Despite this limitation, we refer to the result in Eq. (12) as Floquet-Gibbs limit, because the expression depends only on the quasienergy in the first Brillouin zone, $q_{\alpha}$, and not on the sidebands energies $q_{\alpha}+k\Omega$ with $k\neq 0$.
For larger temperatures, we assume that for all relevant sidebands, approximation (10) holds. Then,
$P_{\alpha}=\frac{1}{2}+\sum_{k}\langle\phi_{\alpha k}|\phi_{\alpha k}\rangle\frac{q_{\alpha}+k\Omega}{4k_{B}T}=f(E_{\alpha})$ (13)
where we have used expression (2) for the mean energy. As we will see below in our numerical example, Eq. (13) predicts the correct behavior for intermediate and large temperatures.
Figure 3: (a) Mean energies and quasienergy splitting on a circle around the cone tip at the first crossing for driving frequency $\Omega=3\Delta$. The angle $\vartheta=0$ corresponds to $\varepsilon=0$ and a driving amplitude $A<A_{0}$, i.e., slightly smaller than the one at the cone tip. (b) Populations for equal coupling to two leads with tunnel rate $\Gamma_{L}=\Gamma_{R}=\Delta/10$ and chemical potential $\mu=0$.
For a numerical confirmation, we focus on the quasienergy crossing shown in Fig. 2(b) and choose the detuning and the amplitude on a circle around the tip with $r=0.1\Delta$ as described at the end of Sec. II. On this circle, the mean energies shown in Fig. 3(a) are interchanged as discussed above. The corresponding populations for a relatively small driving frequency $\Omega=3\Delta$ and both low and intermediate temperature are shown in Fig. 3(b). One can appreciate that the approximation in Eq. (13) fits the numerical result rather well, i.e., a mean-energy state is established. It is also worth mentioning that the continuous interchange of the mean energies and the populations implies that for some value of $\vartheta$, $E_{0}=E_{1}$ and $P_{0}=P_{1}=1/2$. The latter means that the density operator has maximal entropy even at intermediate temperatures. In the absence of driving, such behavior is expected only when the thermal energy by far exceeds all energy splittings. For the driven two-level system in the vicinity of conical intersections, it has been found also for dissipation stemming from an Ohmic heat bath.Kohler (2024)
In the low-temperature limit $k_{B}T=0.01\Delta$, owing to the practically constant quasienergies, Eq. (12) may lead to the conclusion that the populations should be constant on the circle. By contrast, we witness a clear dependence on $\vartheta$, i.e., the Floquet-Gibbs limit is not fulfilled. Also the $\vartheta$-dependence of the zeroth sideband, $\langle\phi_{\alpha 0}|\phi_{\alpha 0}\rangle$, cannot explain the discrepancy.
Figure 4: Population difference as a function of the temperature for detuning $\varepsilon=0$, amplitude $A>A_{0}$ (i.e., $\vartheta=\pi$), and various driving frequencies. The dashed and the dotted lines mark the conjectures of a Floquet-Gibbs limit (FG) and a mean-energy state (ME) in Eqs. (12) and (13), respectively. The mean energy is practically the same for all driving frequencies considered. All other parameters are as in Fig. 3.
To demonstrate that this discrepancy can be attributed to the relatively small driving frequency, we compare the conjectures (12) and (13) with the numerical results for different driving frequencies and temperatures. As quantity of interest, we consider the population difference $\Delta P\equiv|P_{1}-P_{0}|$ for an amplitude slightly above the cone tip, i.e., for $\vartheta=\pi$. The results expected from Eqs. (12) and (13) are $\Delta P=\langle\phi_{\alpha 0}|\phi_{\alpha 0}\rangle\coth(|q_{0}|/2k_{B}T)$ and $\coth(|E_{0}|/2k_{B}T)$, respectively, where we have used $q_{0}+q_{1}=0=E_{0}+E_{1}$. Figure 4 shows that the approximation in Eq. (13) is fulfilled rather well for $k_{B}T\gtrsim\Omega$. However, the Floquet-Gibbs limit (dotted line) only emerges for a relatively large driving frequency $\Omega\ggg\Delta$. In this respect, the present situation is analogous to the one with a bosonic heat bath.Kohler (2024) It is worth mentioning that same conclusions can be drawn from the data for $\vartheta=0$. For this value, however, quasienergies and mean energies have opposite order, $q_{0}<q_{1}$ while $E_{0}>E_{1}$. This leads to a non-monotonic behavior as a function of $k_{B}T$ which is less convenient to analyze.
III.2 Asymmetric wire-lead couplings
Figure 5: (a) The same as Fig. 3(b) but in the absence of the right lead, $\Gamma_{R}=0$, which yields significant deviations from the conjectures in Eqs. (12) and (13). (b) Coefficient $w_{\alpha k}^{(\ell)}=\Gamma_{\ell}|\langle\ell|\phi_{\alpha k}\rangle|^{2}$ appearing in Eq. (9) for the Floquet “ground state” $|\phi_{0}(t)\rangle$ and dot state $|L\rangle$. The dashed lines show for comparison the corresponding coefficient for two leads with equal tunnel rates, $\Gamma_{L}=\Gamma_{R}$.
To highlight the role of equal coupling to both leads, we consider the asymmetric case with the right lead disconnected, $\Gamma_{R}=0$. All other parameters are as before, such that both the quasienergies and the mean energies are unchanged. Nevertheless, the resulting Floquet state populations shown in Fig. 5(a) differ qualitatively from the ones in Fig. 3(b). Even the reflection symmetry at $\vartheta=\pi$ is lost. Obviously, we nor longer have a Floquet-Gibbs-like state or a mean energy state.
The impact of the asymmetry can be understood from the analytic result (9). For a coupling to only the left lead, the projectors $|\ell\rangle\langle\ell|$ in the coefficients $w_{\alpha k}^{(\ell)}$ no longer sum up to unity, such that
$P_{\alpha}\propto\sum_{k}|\langle L|\phi_{\alpha k}\rangle|^{2}f(q_{\alpha}+k\Omega).$ (14)
Figure  5(b) depicts one of the prefactors of the Fermi function in Eq. (14) for the sideband indices $k=0,\pm 1$. It can be appreciated that at $\vartheta=\pi/2$, they are rather small and for some value of $\vartheta$ practically vanish. However, they vanish at different angles. Therefore, their relative difference may be huge, and the symmetry with respect to the sign of $k$ underlying the derivation of Eq. (12) may be significantly broken. As a consequence, a dependence of $P_{\alpha}$ on some sideband energies will remain.
For the high-temperature approximation that leads to the mean-energy state the situation is even clearer. The sideband occupation in the middle of Eq. (13) now contains the projector $|L\rangle\langle L|$. Therefore, the summation no longer yields the mean energy. This is already gradually the case for $\Gamma_{L}\neq\Gamma_{R}$. This reasoning reveals that the symmetry of the dot-lead couplings is a crucial ingredient for a stationary state that can be written as simple function of the quasienergy or the mean energy.
Let finally us remark that the (approximate) localization of the Floquet states for $A=A_{0}$ and small but non-vanishing detuning $\epsilon$ (i.e., $\vartheta=\pi/2$ and $23\pi/$ 2) that explains the data in Fig. 5(b) can be understood form the perturbation theory in Sec. II.1. Since at the tip, the zeroth order of the Floquet Hamiltonian vanishes, the Floquet states are essentially determined by the detuning, which is proportional to $\sigma_{z}$. Hence the natural expectation is a tendency towards localized states.
IV Conclusions
We have addressed the question whether for a driven quantum system coupled to electron reservoirs, Floquet-Gibbs states or mean-energy states emerge. As a basic model, we have considered a two-level system that can be occupied with up to two spinless electrons without taking their interaction into account. This situation allows a good comparison with a former study Kohler (2024) in which dissipation stems from a coupling to a bosonic heat bath. Also here we have focussed on conical intersections, because in their vicinity, quasienergies and mean energies behave rather differently and in a characteristic manner. Then much insight can be gained from the Bloch-Redfield master equation for the single-particle density operator. Within a rotating-wave approximation it can be solved analytically.
A main observation is that no true Floquet-Gibbs states are found, i.e., there seem to be no non-trivial limit in which the populations match the Fermi function evaluated at the quasienergies. Nevertheless, there are situations in which the deviation is determined by the quasienergy in the first Brillouin zone, such that one may speak of Floquet-Gibbs-like states. As for the dissipative two-level system, this requires rather large driving frequencies. For smaller frequencies and intermediate temperatures, the stationary populations are directly given by a Fermi function with the mean energy of the Floquet states as argument. Thus, also here the mean-energy states seem rather generic, at least under the condition of a sufficiently symmetric setup. In particular, the tunnel couplings of the sites to the respective leads have to be equal. For larger systems, for example for models that describe conducting molecules, such equal coupling of all sites to an electron reservoir may be essential for the emergence of Floquet-Gibbs-like states or mean-energy states.
Acknowledgements.
This work was supported by the Spanish Ministry of Science, Innovation, and Universities (Grant No. PID2023-149072NB-I00), and by the CSIC Research Platform on Quantum Technologies PTI-001.
Data Availability Statement
The numerical data shown in the figures is available from the author upon reasonable request.
Appendix A Floquet Hamiltonian near the cone tip
For a perturbation theory in the vicinity of the cone tip, we follow Ref. Kohler, 2024 and separate the Floquet Hamiltonian into two parts, namely the zeroth-order contribution
$\mathcal{H}_{0}=\frac{\Delta}{2}\sigma_{x}+\frac{A_{0}}{2}\sigma_{z}\cos(\Omega t)-i\frac{\partial}{\partial t}\equiv H_{0}(t)-i\frac{\partial}{\partial t},$ (15)
with $A_{0}$ assuming a value for which the quasienergies with $q=0$ are degenerate. Since $\mathcal{H}_{0}$ obeys generalized parity, $G\mathcal{H}_{0}G^{-1}=\mathcal{H}_{0}$, its two non-equivalent Floquet states can be classified as even or odd. They will be denoted by $|\varphi_{\pm}(t)\rangle$ with mean energies $E_{\pm}$. The perturbation
$\mathcal{H}_{1}=\frac{\varepsilon}{2}\sigma_{z}+\frac{A-A_{0}}{2}\sigma_{z}\cos(\Omega t)$ (16)
will be treated with degenerate perturbation theory. To do so, we calculate the matrix elements of $\mathcal{H}_{1}$. Since $G\sigma_{z}G^{-1}=-\sigma_{z}$ and $G\sigma_{z}\cos(\Omega t)G^{-1}=\sigma_{z}\cos(\Omega t)$, in the new basis the diagonal matrix elements of the first term in Eq. (16) and the off-diagonal elements of the second term vanish.
The remaining matrix elements of $\sigma_{z}$ can be evaluated straightforwardly to yield $b/2$ with $b$ given in Eq. (5). The diagonal elements of the time-dependent term can be obtain upon noticing that
$\sigma_{z}\cos(\Omega t)=\big{(}2H_{0}(t)-\Delta\sigma_{x}\big{)}/A_{0}.$ (17)
Together with the relation $E_{-}=\langle\!\langle\varphi_{-}|H_{0}(t)|\varphi_{-}\rangle\!\rangle=-E_{+}$ follows the first term of the effective Floquet Hamiltonian (3) with the coefficient $a$ in Eq. (4).